a chessboard has an area of 324 square inches. there is a 1 in border around the 64 squares on the board. what is the length of one side of the region containing the small squares?
How would you figure out the answer

Answer by scott8148(6628)

the length of the whole board is 2 inches more than the length of the squares part (one inch all around)

(L + 2)^2 = 324 ___ L^2 + 4L + 4 = 324 ___ L^2 + 4L - 320 = 0

factoring ___ (L + 20)(L - 16) = 0

L + 20 = 0 ___ L = - 20 ___ negative value not realistic

L - 16 = 0 ___ L = 16